# Structured Kolmogorov-Arnold Neural ODEs for Interpretable Learning and Symbolic Discovery of Nonlinear Dynamics

### v1

@@ Abstract
Understanding and modeling nonlinear dynamical systems is a fundamental problem across scientific and engineering domains. While deep learning has demonstrated remarkable potential for learning complex system behavior, achieving models that are both highly accurate and physically interpretable remains a major challenge. To address this, we propose Structured Kolmogorov-Arnold Neural ODEs (SKANODEs), a novel framework that integrates structured state-space modeling with the Kolmogorov-Arnold Network (KAN). SKANODE first employs a fully trainable KAN as a universal function approximator within a structured Neural ODE framework to perform virtual sensing, recovering latent states that correspond to physically interpretable quantities such as positions and velocities. Once this structured latent representation is established, we exploit the symbolic regression capability of KAN to extract compact and interpretable expressions for the system’s governing dynamics. The resulting symbolic expression is then substituted back into the Neural ODE framework and further calibrated through continued training to refine its coefficients, enhancing both the precision of the discovered equations and the predictive accuracy of system responses. Extensive experiments on both simulated and real-world systems demonstrate that SKANODE achieves superior performance while offering interpretable, physics-consistent models that uncover the underlying mechanisms of nonlinear dynamical systems.
@@ 1 Introduction
To overcome these limitations, we propose Structured Kolmogorov-Arnold Neural ODEs (SKANODEs), a novel framework that unifies continuous-time neural modeling, structured physical inductive biases, and end-to-end symbolic discovery of governing equations within a single differentiable pipeline. At the core of SKANODE is the Kolmogorov-Arnold Network (KAN) [22], a recent neural architecture that supports smooth transitions between black-box function approximation and symbolic representation learning. KAN employs adaptive, spline-based basis functions that are differentiable and expressive, while retaining the structure necessary for symbolic interpretability.
@@
Contributions: We propose Structured Kolmogorov-Arnold Neural ODEs (SKANODEs), a new modeling framework for interpretable and data-driven equation discovery of nonlinear dynamical systems, under partial observability. Our contributions include:
@@
    A physically structured NODE architecture, incorporating both a structured state-space and a structured observation model. This enables interpretable latent state learning from indirect sensor measurements (e.g., accelerations), performing virtual sensing without requiring full observation of physical coordinates.
@@
Through extensive evaluations on both simulated and real-world systems, we demonstrate that SKANODE achieves significant predictive accuracy while enabling transparent equation-level understanding of nonlinear dynamics. This opens up new possibilities for applying deep learning in scientific modeling, engineering diagnostics, and physics-based decision making.
@@ 2.3 Symbolic Equation Discovery and Kolmogorov–Arnold Networks
To improve scalability and expressiveness, differentiable symbolic regression frameworks such as AI Feynman [49], PySR [50], and neural symbolic regression (e.g., [19, 51, 52]) employ neural-guided search or evolutionary optimization to discover analytic expressions. These methods offer greater automation but retain critical limitations: they often function as post hoc pipelines requiring clean, fully observed data and are rarely integrated directly into architectures like NODEs.
@@
To overcome these limitations, we adopt the Kolmogorov–Arnold Network (KAN)—a neural architecture grounded in the Kolmogorov–Arnold representation theorem and introduced in recent literature [22]. KAN replaces traditional MLPs with networks whose activation function nodes are parameterized as learnable spline functions, enabling both:
@@ 2.4 State-Space Representation
It is reminded that we here focus on second-order dynamics because they are prevalent in vibrational analysis, where the system states can be interpreted as displacement and velocity. These dynamics are significant in various engineering and physical systems, such as mechanical oscillators, electrical circuits, and structural vibrations [60]. This approach also serves as a foundation for extending the analysis to more complex, higher-order systems. For completeness, we remind that the state-space form naturally applies to higher-order dynamics. Given a general nonlinear $n$-order ODE of the form:
@@
we define the state variable $\mathbf{z}=[z_{1},z_{2},z_{3},...,z_{n}]$ as:
@@ 3.1 Structured State-Space Modeling
To address this, we introduce a structured state-space formulation tailored for modeling such systems using Neural ODEs. We define the latent state as:
@@
where $\mathbf{u}(t)\in\mathbb{R}^{p}$ is the external input (e.g., force), and $\mathbf{y}(t)\in\mathbb{R}^{d}$ are the measured accelerations. The same function $h_{\theta}$ governs both the transition dynamics and the observation process, ensuring consistency between the latent state evolution and observable measurements. We denote by $\mathcal{H}$ the function space of candidate functions for the governing dynamics, consisting of all functions $h:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{d}$ parameterized by the trainable network $h_{\theta}$. In the SKANODE framework, $\mathcal{H}$ is realized either by fully trainable Kolmogorov–Arnold Networks (KAN) during the universal approximation stage, or by their extracted symbolic representations after symbolic equation discovery.
@@
Training Objective. To estimate the model parameters $\theta$, we define the loss function directly as the discrepancy between the predicted and measured accelerations:
@@ 3.2 KAN-Based Two-Stage Learning
To facilitate interpretable equation discovery within the structured state-space Neural ODE framework, we incorporate the Kolmogorov–Arnold Network (KAN), a spline-based architecture capable of acting both as a universal function approximator and a symbolic expression learner.
@@
In the first stage, we model the acceleration dynamics using $\text{KAN}_{\text{approx}}$, a fully trainable Kolmogorov–Arnold Network configured as a universal approximator. This network models the acceleration dynamics based on the structured Neural ODE formulation:
@@
Once $\text{KAN}_{\text{approx}}$ has been trained and meaningful latent states are recovered, we use these learned displacement and velocity signals as inputs to a second KAN, denoted $\text{KAN}_{\text{symbolic}}$. This KAN is utilized in symbolic mode, enabling automatic extraction of analytic expressions for the system’s governing dynamics.
@@ 3.3 Properties of SKANODE
We now state a formal identifiability result establishing that under certain conditions SKANODE can recover the true governing dynamics of the system.
@@ 4 Results
To demonstrate the enhanced interpretability and precision of the dynamics learned by our proposed method, we assess SKANODE through a series of examples involving both simulated and real-world nonlinear dynamical systems. We first apply SKANODE to two classic benchmark systems broadly used in nonlinear dynamics research: the Duffing oscillator and the Van der Pol oscillator [60]. These synthetic examples illustrate SKANODE’s ability to accurately recover physically meaningful latent states and to extract precise symbolic expressions for the governing equations. We then extend the evaluation to real-world data by analyzing ground vibration measurements from an F-16 aircraft, aiming to model the nonlinear interactions at the interface between the aircraft’s right wing and its payload. The results across these experiments further support the theoretical properties discussed in the previous section, demonstrating the effectiveness of SKANODE in practical applications.
For all comparative experiments, care was taken to ensure a consistent training setup across models. The only architectural difference lies in the use of the KAN in SKANODE, whereas ANODE and SONODE adopt MLPs. The architectures of the KAN modules used in SKANODE are visualized in the corresponding experimental subsections. For MLP-based models, a 5-layer network with 32 hidden units per layer is adopted throughout. Identical data preprocessing, batch sizes, learning rates schedules, and total training epochs were applied across all models to ensure a fair comparison. The data and code used in this paper are publicly available on GitHub at https://github.com/liouvill/SKANODE.
@@ 4.1 Duffing Oscillator
Figure 2: Results on the Duffing oscillator. Top left: Symbolic governing equation discovered by $\text{KAN}_{\text{symbolic}}$; a node receiving displacement input $x$ clearly exhibits the expected cubic nonlinear relationship. Top right: Predicted system observables (accelerations) using the SKANODE framework. SKANODE achieves high accuracy in reconstructing accelerations, directly from latent dynamics without requiring a separate observation model. Bottom: Inferred latent state variables compared to ground truth. The latent states recovered by SKANODE correspond directly to physically meaningful displacement and velocity trajectories, in contrast to ANODE and SONODE, whose latent states lack clear physical interpretation and serve primarily as abstract representations for reconstructing observables.
We evaluate and compare the performance of SKANODE against the equivalent state of the art in the field, namely ANODE and SONODE, on this system. The symbolic governing equation extracted by $\text{KAN}_{\text{symbolic}}$ within SKANODE is shown in Figure 2, where the identified expression accurately captures the true system dynamics. Notably, the extracted symbolic network exhibits a distinct cubic node connected to the displacement input $x$, consistent with the characteristic cubic stiffness nonlinearity inherent in the Duffing oscillator. The predicted accelerations and recovered latent state trajectories are also presented. While all three models achieve reasonable accuracy in reconstructing the observable accelerations, only SKANODE successfully recovers latent states that correspond directly to physically meaningful displacement and velocity trajectories. This advantage stems from its structured state-space design and the symbolic learning capability of the Kolmogorov–Arnold Network.
@@
Noise level Model acceleration (measurement) displacement velocity
ANODE (3.15 $\pm$ 1.42)×$\times 10^{-2}$ (9.95 $\pm$ 0.32)×$\times 10^{-1}$ 6.06 $\pm$ 5.17
0 SONODE (1.52 $\pm$ 0.71)×$\times 10^{-2}$ 1.01 $\pm$ 0.02 9.46 $\pm$ 0.14
S3NODE (1.11 $\pm$ 0.51)×$\times\mathbf{10^{-3}}$ (1.45 $\pm$ 0.76)×$\times\mathbf{10^{-3}}$ (4.53 $\pm$ 2.78)×$\times\mathbf{10^{-4}}$
SKANODE (3.29 $\pm$ 0.01)×$\times\mathbf{10^{-5}}$ (2.81 $\pm$ 0.01)×$\times\mathbf{10^{-4}}$ (1.31 $\pm$ 0.01)×$\times\mathbf{10^{-5}}$
@@
0.001 SONODE (5.41 $\pm$ 6.13)×$\times 10^{-2}$ 1.01 $\pm$ 0.02 9.33 $\pm$ 0.33
S3NODE (6.78 $\pm$ 3.43)×$\times\mathbf{10^{-4}}$ (1.08 $\pm$ 0.49)×$\times\mathbf{10^{-3}}$ (2.85 $\pm$ 1.84)×$\times\mathbf{10^{-4}}$
SKANODE (3.36 $\pm$ 0.01)×$\times\mathbf{10^{-5}}$ (2.80 $\pm$ 0.01)×$\times\mathbf{10^{-4}}$ (1.30 $\pm$ 0.01)×$\times\mathbf{10^{-5}}$
@@
0.005 SONODE (2.19 $\pm$ 2.86)×$\times 10^{-1}$ (7.92 $\pm$ 3.07)×$\times 10^{-1}$ 6.89 $\pm$ 3.59
S3NODE (1.25 $\pm$ 0.99)×$\times\mathbf{10^{-3}}$ (1.46 $\pm$ 1.07)×$\times\mathbf{10^{-3}}$ (5.09 $\pm$ 4.41)×$\times\mathbf{10^{-4}}$
SKANODE (5.90 $\pm$ 0.01)×$\times\mathbf{10^{-5}}$ (2.90 $\pm$ 0.01)×$\times\mathbf{10^{-4}}$ (1.40 $\pm$ 0.01)×$\times\mathbf{10^{-5}}$
@@
0.01 SONODE (1.76 $\pm$ 0.57)×$\times 10^{-2}$ 1.01 $\pm$ 0.02 9.39 $\pm$ 0.22
S3NODE (1.18 $\pm$ 1.00)×$\times\mathbf{10^{-3}}$ (4.64 $\pm$ 1.97)×$\times\mathbf{10^{-4}}$ (4.07 $\pm$ 4.62)×$\times\mathbf{10^{-4}}$
SKANODE (1.38 $\pm$ 0.01)×$\times\mathbf{10^{-4}}$ (2.93 $\pm$ 0.02)×$\times\mathbf{10^{-4}}$ (1.44 $\pm$ 0.01)×$\times\mathbf{10^{-5}}$
@@
To quantitatively assess model performance, we compute mean squared errors (MSE) for both predicted accelerations and inferred latent states compared to ground truth. The mean and standard deviation of MSE values across three independent runs are reported in Table 1. All models achieve similar performance in predicting accelerations, though SKANODE consistently shows slightly superior accuracy. However, significant differences arise in the recovery of latent states: SKANODE achieves the lowest MSE and variance, demonstrating its ability to extract physically interpretable dynamics reliably across repeated trials. In contrast, ANODE exhibits high variance and poor latent state interpretability, verifying that ANODE tends to learn arbitrary latent representations to reconstruct observables. SONODE yields more stable but still non-interpretable latent trajectories, as its structure lacks explicit alignment with the true physical coordinates. Furthermore, we also evaluate a structured state-space Neural ODE using a conventional multilayer perceptron (MLP) in place of KAN—referred to as S3NODE in Table 1—implemented as a 5-layer MLP with 32 hidden units per layer. The results demonstrate that the superior performance of SKANODE is not only due to its structural inductive biases in the latent and observation models, but also to the symbolic discovery capability of KAN, which enhances both interpretability and predictive accuracy, by uncovering explicit governing equations and improving the estimation of system quantities.
@@
Importantly, SKANODE not only improves interpretability but also contributes to better predictive performance, highlighting the value of incorporating physics-informed structure and symbolic discovery. To further examine robustness, Table 1 presents the results under varying levels of observation noise. Notably, ANODE encounters severe convergence issues under noisy conditions, while SKANODE maintains stable performance across all scenarios.

### v2

@@ Abstract
Understanding and modeling nonlinear dynamical systems is a fundamental challenge across science and engineering. Deep learning has shown remarkable potential for capturing complex system behavior, yet achieving models that are both accurate and physically interpretable remains difficult. To address this, we propose Structured Kolmogorov-Arnold Neural ODEs (SKANODEs), a framework that integrates structured state-space modeling with Kolmogorov-Arnold Networks (KANs). Within a Neural ODE architecture, SKANODE employs a fully trainable KAN as a universal function approximator to perform virtual sensing, recovering latent states that correspond to interpretable physical quantities such as displacements and velocities. Leveraging KAN’s symbolic regression capability, SKANODE then extracts compact, interpretable expressions for the system’s governing dynamics. Extensive experiments on simulated and real-world systems demonstrate that SKANODE achieves superior predictive accuracy, discovers physics-consistent dynamics, and reveals complex nonlinear behavior. Notably, it identifies hysteretic behavior in an F-16 aircraft and recovers a concise symbolic equation describing this phenomenon. SKANODE thus enables interpretable, data-driven discovery of physically grounded models for complex nonlinear dynamical systems.
@@ 1 Introduction
To overcome these limitations, a novel framework, termed Structured Kolmogorov-Arnold Neural ODEs (SKANODEs), is proposed, which unifies continuous-time neural modeling, structured physical inductive biases, and end-to-end symbolic discovery of governing equations within a single differentiable pipeline. At the core of SKANODE is the Kolmogorov-Arnold Network (KAN) [22], a recent neural architecture that supports smooth transitions between black-box function approximation and symbolic representation learning. KAN employs adaptive, spline-based basis functions that are differentiable and expressive, while retaining the structure necessary for symbolic interpretability.
@@
Contributions: A new modeling framework, termed Structured Kolmogorov-Arnold Neural ODEs (SKANODEs), is proposed for interpretable and data-driven equation discovery of nonlinear dynamical systems, under partial observability. The main contributions of this work include:
@@
    A physically structured Neural ODE architecture, incorporating both a structured state-space and a structured observation model. This enables interpretable latent state learning from indirect sensor measurements (e.g., accelerations), performing virtual sensing without requiring full observation of physical coordinates.
@@
  • 4.
    Extensive experimental validation on both simulated and real-world systems, including the canonical Duffing and Van der Pol oscillators, where SKANODE successfully recovers governing equations consistent with ground truth, and a real-world F-16 aircraft system, where the model identifies hysteretic behavior and derives an interpretable symbolic equation capturing this nonlinear phenomenon.
Through extensive evaluations on both simulated and real-world systems, it is demonstrated that SKANODE achieves significant predictive accuracy while enabling transparent equation-level understanding of nonlinear dynamics. This opens up new possibilities for applying deep learning in scientific modeling, engineering diagnostics, and physics-based decision making.
@@ 2.3 Symbolic Equation Discovery and Kolmogorov–Arnold Networks
To improve scalability and expressiveness, differentiable symbolic regression frameworks such as AI Feynman [49], PySR [50], and neural symbolic regression (e.g., [19, 51, 52]) employ neural-guided search or evolutionary optimization to discover analytic expressions. These methods offer greater automation but retain critical limitations: they often function as post hoc pipelines requiring clean, fully observed data and are rarely integrated directly into architectures like NODEs [48].
@@
To overcome these limitations, the Kolmogorov–Arnold Network (KAN) is adopted—a neural architecture grounded in the Kolmogorov–Arnold representation theorem and introduced in recent literature [22]. KAN replaces traditional MLPs with networks whose activation function nodes are parameterized as learnable spline functions, enabling both:
@@
Sparsification. To support interpretability and generalization, KAN also incorporates a built-in sparsification mechanism. Unlike MLPs, which typically impose sparsity on linear weights, KAN applies L1 regularization directly on the learnable activation functions. Specifically, the L1 norm of an activation function $\phi$ is defined as the average magnitude of its output over a batch of $N_{p}$ inputs $\{x^{(s)}\}_{s=1}^{N_{p}}$:
$|\phi|_{1}\equiv\frac{1}{N_{p}}\sum_{s=1}^{N_{p}}\left|\phi(x^{(s)})\right|.$ (4)
For a KAN layer $\mathbf{\Phi}$ with $n_{\text{in}}$ inputs and $n_{\text{out}}$ outputs, we denote $\phi_{i,j}$ as the learnable activation function connecting input $i$ to output $j$. The L1 norm of the entire layer is the sum of L1 norms of all activation functions:
$|\mathbf{\Phi}|_{1}\equiv\sum_{i=1}^{n_{\text{in}}}\sum_{j=1}^{n_{\text{out}}}|\phi_{i,j}|_{1}.$ (5)
To further enhance sparsity, KAN introduces an entropy regularization term that encourages selective activation. The entropy of a layer is defined as:
$S(\mathbf{\Phi})\equiv-\sum_{i=1}^{n_{\text{in}}}\sum_{j=1}^{n_{\text{out}}}\frac{|\phi_{i,j}|_{1}}{|\mathbf{\Phi}|_{1}}\log\left(\frac{|\phi_{i,j}|_{1}}{|\mathbf{\Phi}|_{1}}\right),$ (6)
which penalizes a uniform distribution over activation magnitudes and encourages a sparse subset of nodes to dominate.
The total training objective combines the prediction loss with both L1 and entropy regularization across all KAN layers indexed by $l=0,\dots,L-1$:
$\ell_{\text{total}}=\ell_{\text{pred}}+\lambda\left(\mu_{1}\sum_{l=0}^{L-1}|\mathbf{\Phi}^{l}|_{1}+\mu_{2}\sum_{l=0}^{L-1}S(\mathbf{\Phi}^{l})\right),$ (7)
where $\lambda$ controls the overall regularization strength, and $\mu_{1},\mu_{2}$ (typically set to 1) balance the two regularization components.
This sparsification framework allows KAN to learn parsimonious and interpretable structures while preserving representational capacity. In this work, this built-in regularization strategy is adopted as part of the proposed symbolic modeling framework.
@@ 2.4 State-Space Representation
It is reminded that we here focus on second-order dynamics because they are prevalent in vibrational analysis, where the system states can be interpreted as displacement and velocity. These dynamics are significant in various engineering and physical systems, such as mechanical oscillators, electrical circuits, and structural vibrations [60]. This approach also serves as a foundation for extending the analysis to more complex, higher-order systems. For completeness, it is noted that the state-space form naturally applies to higher-order dynamics. Given a general nonlinear $n$-order ODE of the form:
@@
the state variable $\mathbf{z}=[z_{1},z_{2},z_{3},...,z_{n}]$ is defined as:
@@ 3.1 Structured State-Space Modeling
To address this, a structured state-space formulation is introduced, tailored for modeling such systems using Neural ODEs. The latent state is defined as:
@@
where $\mathbf{u}(t)\in\mathbb{R}^{p}$ is the external input (e.g., force), and $\mathbf{y}(t)\in\mathbb{R}^{d}$ are the measured accelerations. The same function $h_{\theta}$ governs both the transition dynamics and the observation process, ensuring consistency between the latent state evolution and observable measurements. Let $\mathcal{H}$ denote the function space of candidate functions for the governing dynamics, consisting of all functions $h:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{d}$ parameterized by the trainable network $h_{\theta}$. In the SKANODE framework, $\mathcal{H}$ is realized either by fully trainable Kolmogorov–Arnold Networks (KAN) during the universal approximation stage, or by their extracted symbolic representations after symbolic equation discovery.
@@
Training Objective. To estimate the model parameters $\theta$, the loss function is defined as the discrepancy between the predicted and measured accelerations:
@@
To avoid overfitting, the sparsification techniques detailed in Section 2.3 are incorporated into the training procedure. These include L1 regularization and entropy regularization, which control the model complexity and improve generalization. Additionally, in $\text{KAN}_{\text{symbolic}}$, several linear nodes are predefined for specific terms, such as the linear components of the damping ratio and stiffness. These linear relationships are fixed apriori, with only the affine weights remaining trainable. Nonlinear terms, such as the cubic term in the Duffing oscillator and the quadratic term in the Van der Pol oscillator, are fully discovered by the network, including both the functional form and the affine parameters.
@@ 3.2 KAN-Based Two-Stage Learning
To facilitate interpretable equation discovery within the structured state-space Neural ODE framework, the Kolmogorov–Arnold Network (KAN) is incorporated, which is a spline-based architecture capable of acting both as a universal function approximator and a symbolic expression learner.
@@
In the first stage, the acceleration dynamics are modeled using $\text{KAN}_{\text{approx}}$, a fully trainable Kolmogorov–Arnold Network configured as a universal approximator. This network models the acceleration dynamics based on the structured Neural ODE formulation:
@@
Once $\text{KAN}_{\text{approx}}$ has been trained and meaningful latent states are recovered, the learned displacement and velocity signals are used as inputs to a second KAN, denoted $\text{KAN}_{\text{symbolic}}$. This KAN is utilized in symbolic mode, enabling automatic extraction of analytic expressions for the system’s governing dynamics.
@@ 3.3 Properties of SKANODE
A formal identifiability result is now stated, establishing that under certain conditions SKANODE can recover the true governing dynamics of the system.
@@ 4 Results
To demonstrate the enhanced interpretability and precision of the dynamics learned by the proposed method, SKANODE is assessed through a series of examples involving both simulated and real-world nonlinear dynamical systems. SKANODE is first applied to two classic benchmark systems broadly used in nonlinear dynamics research: the Duffing oscillator and the Van der Pol oscillator [60]. These synthetic examples illustrate SKANODE’s ability to accurately recover physically meaningful latent states and to extract precise symbolic expressions for the governing equations. The evaluation is then extended to real-world data by analyzing ground vibration measurements from an F-16 aircraft, aiming to model the nonlinear interactions at the interface between the aircraft’s right wing and its payload. The results across these experiments further support the theoretical properties discussed in the previous section, demonstrating the effectiveness of SKANODE in practical applications.
For all comparative experiments, care was taken to ensure a consistent training setup across models. The only architectural difference lies in the use of the KAN in SKANODE, whereas ANODE and SONODE adopt MLPs. The architectures of the KAN modules used in SKANODE are visualized in the corresponding experimental subsections. In our implementation, B-splines are adopted as the spline basis, with 5th-order B-splines specifically employed. The activation functions are modeled via B-splines with a residual connection, which is a typical design feature of KANs. Each KAN is initialized with 6 knots (forming 5 intervals), and the knots are adaptively updated during the training process. For MLP-based models, a 5-layer network with 32 hidden units per layer is adopted throughout. Identical data preprocessing, batch sizes, learning rates schedules, and total training epochs were applied across all models to ensure a fair comparison. The experiments are conducted on a local workstation with an NVIDIA GeForce GTX 5080 GPU. Each experiment is run for 2,000 epochs. The data and code used in this paper are publicly available on GitHub at https://github.com/liouvill/SKANODE.
@@ 4.1 Duffing Oscillator
Figure 2: Results on the Duffing oscillator. Top left: Symbolic governing equation discovered by $\text{KAN}_{\text{symbolic}}$, compared against the numerical baseline. The identified node receiving displacement input $x$ correctly captures the expected cubic nonlinearity. Top right: Predicted system observables (accelerations) obtained using the SKANODE framework. SKANODE accurately reconstructs accelerations directly from the inferred latent dynamics, without the need for a separate observation model. Bottom: Inferred latent state variables compared to ground truth. The latent states recovered by SKANODE correspond closely to physically meaningful displacement and velocity trajectories, unlike ANODE and SONODE, whose latent states remain abstract representations without clear physical interpretation.
The performance of SKANODE is evaluated and compared against the equivalent state of the art in the field, namely ANODE and SONODE, on this system, as shown in the top-right subfigure of Figure 2. For benchmarking the symbolic learning capability, a numerical baseline is constructed by directly integrating the measured acceleration signals to obtain velocity and displacement, which are then supplied to SINDy to extract a compact symbolic expression. The predicted trajectories from this numerical benchmark are generated by solving the discovered equations and serve to assess the accuracy of the identified dynamics. It is important to note that this numerical baseline serves solely as a reference for benchmarking the symbolic discovery process; it is not a predictor by itself, as it merely integrates the measured signals without learning the underlying dynamics prior to extracting a symbolic form of the governing equations, and its predictions completely depend on the quality of the extracted equation. In contrast, the proposed SKANODE framework functions both as a dynamics learner—capable of end-to-end prediction of system responses—and as a symbolic extractor that uncovers governing equations from the learned latent quantities.
The symbolic governing equation extracted by $\text{KAN}_{\text{symbolic}}$ within SKANODE is shown in Figure 2, where the identified expression accurately captures the true system dynamics. Notably, the extracted symbolic network exhibits a distinct cubic node connected to the displacement input $x$, consistent with the characteristic cubic stiffness nonlinearity inherent in the Duffing oscillator. The predicted accelerations and recovered latent state trajectories are also presented. While all three models achieve reasonable accuracy in reconstructing the observable accelerations, only SKANODE successfully recovers latent states that correspond directly to physically meaningful displacement and velocity trajectories. This advantage stems from its structured state-space design and the symbolic learning capability of the Kolmogorov–Arnold Network.
@@
Noise
level
Model
acceleration
displacement velocity
[…]
@@
[…]
@@
[…]
@@
[…]
@@
To quantitatively evaluate model performance, mean squared errors (MSE) are computed for both predicted accelerations and inferred latent states against the ground truth. The mean and standard deviation of MSE values over three independent runs are reported in Table 1. All models achieve similar performance in predicting accelerations, while SKANODE consistently shows slightly superior accuracy. More pronounced differences appear in the recovery of latent states: SKANODE achieves the lowest MSE and variance, indicating its ability to extract physically interpretable dynamics reliably across repeated trials. In contrast, ANODE displays high variance and poor latent state interpretability, confirming its tendency to learn arbitrary latent representations. SONODE produces more stable but still non-interpretable latent trajectories, as its structure does not explicitly align with the true physical coordinates. Furthermore, a structured state-space Neural ODE using a conventional multilayer perceptron (MLP) in place of KAN is also evaluated—referred to as S3NODE in Table 1—implemented as a 5-layer MLP with 32 hidden units per layer. These results indicate that the advantage of SKANODE stems both from its structured inductive biases in the latent and observation models and from the symbolic discovery capability of KAN, which improves interpretability and predictive accuracy by uncovering governing equations and enhancing the estimation of system quantities.
Table 2: Model performance measured in SSIM for Duffing oscillator experiments subjected to different levels of observation noise. Structural Similarity Index (SSIM) is used to quantify the perceptual similarity between predicted and ground-truth signals. The observational noise is assumed to be Gaussian, randomly sampled from the distribution $\mathcal{N}(0,\sigma^{2})$, and the value listed in the noise level column indicates the value of standard deviation $\sigma$. ANODE fails to converge even at the minimal noise level of 0.001.
Noise
level
Model
acceleration
displacement velocity
Numerical 0.9759 0.5822 0.9317
ANODE 0.8612 $\pm$ 0.0471 0.0335 $\pm$ 0.0010 0.0064 $\pm$ 0.0013
0 SONODE 0.9240 $\pm$ 0.0253 0.0368 $\pm$ 0.0013 0.0365 $\pm$ 0.0016
S3NODE 0.9850 $\pm$ 0.0083 0.6990 $\pm$ 0.0847 0.9623 $\pm$ 0.0209
SKANODE 0.9997 $\pm$ 0.0000 0.8689 $\pm$ 0.0000 0.9974 $\pm$ 0.0000
Numerical 0.9873 0.7285 0.9599
ANODE NaN NaN NaN
0.001 SONODE 0.8623 $\pm$ 0.1085 0.0366 $\pm$ 0.0009 0.0370 $\pm$ 0.0024
S3NODE 0.9895 $\pm$ 0.0056 0.7394 $\pm$ 0.0681 0.9720 $\pm$ 0.0149
SKANODE 0.9997 $\pm$ 0.0000 0.8691 $\pm$ 0.0000 0.9974 $\pm$ 0.0000
Numerical 0.9710 0.4690 0.9173
ANODE NaN NaN NaN
0.005 SONODE 0.6693 $\pm$ 0.3518 0.0309 $\pm$ 0.0093 0.0406 $\pm$ 0.0053
S3NODE 0.9845 $\pm$ 0.0101 0.7171 $\pm$ 0.0969 0.9605 $\pm$ 0.0272
SKANODE 0.9995 $\pm$ 0.0000 0.8663 $\pm$ 0.0000 0.9972 $\pm$ 0.0000
Numerical 0.8667 0.2673 0.6864
ANODE NaN NaN NaN
0.01 SONODE 0.9129 $\pm$ 0.0218 0.0374 $\pm$ 0.0015 0.0355 $\pm$ 0.0009
S3NODE 0.9865 $\pm$ 0.0120 0.8452 $\pm$ 0.0458 0.9692 $\pm$ 0.0291
SKANODE 0.9990 $\pm$ 0.0000 0.8654 $\pm$ 0.0000 0.9971 $\pm$ 0.0000
@@
Importantly, SKANODE not only improves interpretability but also contributes to better predictive performance, underscoring the value of incorporating physics-encoded structure and symbolic discovery capability. To further assess robustness, Table 1 presents results under varying levels of observation noise. It is noted that ANODE experiences severe convergence issues under noisy conditions, even at the minimal noise level, whereas SKANODE maintains stable performance across all scenarios.
To complement MSE and provide a more diagnostic perspective, the Structural Similarity Index (SSIM) is also reported. SSIM captures the perceptual similarity between predicted and ground-truth trajectories. Results across different levels of observation noise are summarized in Table 2. The SSIM analysis reinforces the conclusions drawn from MSE: SKANODE consistently attains near-perfect scores (>0.99>0.99 across all states and noise levels), indicating that its predictions are not only numerically accurate but also structurally faithful. S3NODE also demonstrates strong performance but remains consistently below SKANODE, while SONODE yields moderate similarity and ANODE fails to converge under noisy conditions.
Taken together, these results confirm that the superiority of SKANODE arises from both its structured inductive biases and symbolic discovery capability, which jointly enhance interpretability, predictive accuracy, and robustness compared to conventional neural ODE variants.